You are considering two options: making your own product or buying the product from a supplier. If you make the product, you will need a machine that would cost you $80,000. You can sell this machine after 5 years by $5,000. Other operating and maintenance cost were calculated to be $20 per unit. The supplier sell the product for $30 per unit. The demand of this product per year is 200,000 units. If the MARR for you company is 10%. What option should you select?

Decision can not be made

Buy

Make

Indifferent
Answer #1

Let the breakeven quantity be X, then

Annual cost of making the product = 80000 *(A/P,10%,5) + 20 * X - 5000 * (A/F,10%,5)

= 80000 *0.263797 + 20 * X - 5000 *0.163797

= 20284.81 + 20 * X

Annual cost of buying the product = 30 * X

As per given condition

30 *X = 20284.81 + 20 * X

10 * X = 20284.81

X = 2028.48 ~ 2029 (Next whole number)

As demand is more than breakeven qty, company should make the product